OpenStudy (anonymous):

Factor the expression 5X²-22X-15

OpenStudy (anonymous):

Is there a GCF in this expression?

OpenStudy (anonymous):

5x^2-25x+3x-15=0 5x(x-5)+3(x-5)=0 (5x+3)(x-5)=0 x=5,-3/5